OpenStudy (calculusxy):

Solve using elimination. 0 = -5y + 36 + 9x 18 = -6y + 4x

OpenStudy (calculusxy):

@Michele_Laino

OpenStudy (anonymous):

Did you mean for the first equation to have 2 x terms?

OpenStudy (calculusxy):

Sorry I just changed it now.

OpenStudy (anonymous):

Ahh ok, it just seemed odd for an elimination problem.

OpenStudy (michele_laino):

for example, we can multiply both sides of first equation by 6 and we can multiply both sides of the second equation by -5 so we get this: \[\huge \begin{gathered} 0 = - 30y + 216 + 54x \hfill \\ \hfill \\ - 90 = 30y - 20x \hfill \\ \end{gathered} \] now, please add together such equations

OpenStudy (calculusxy):

ok

OpenStudy (michele_laino):

hint: left side: \[\huge0 - 90 = ...?\]

OpenStudy (michele_laino):

whereas right side is: \[\huge 30y + 216 + 54x + 30y - 20x = ...?\]

OpenStudy (michele_laino):

oops.. I made a typo: right side: \[ \Large - 30y + 216 + 54x + 30y - 20x = ...?\]

OpenStudy (calculusxy):

What I did first is i had the equation (-36 = 5y + 9x)6 (18 = -6y + 45)-5 -216 = 30y + 54x -90 = 30y - 20x -306 = 34x

OpenStudy (calculusxy):

i know i made a mistake in here

OpenStudy (michele_laino):

it is right!

OpenStudy (michele_laino):

now, please solve for \(x\)

OpenStudy (michele_laino):

what is \(-306/34=...?\)

OpenStudy (calculusxy):

-9?

OpenStudy (michele_laino):

that's right, now please replace \(x=-9\) into the first equaton, and solve for \(y\)